OpenStudy (anonymous):

slove equations and put it into a matrices 4t-5r=21 3t-7r=18

OpenStudy (lilg132):

4t - 5r = 21 (multiply all by 7) 3t - 7r = 18 (multiply all by -5) 28t - 35r = 147 -15t + 35r = 90 add the two equations together 13t =237 t = 18.230769230769230769230769230769 4(18.230769230769230769230769230769) - 5r = 21 72.923076923076923076923076923077 - 5r = 21 72.923076923076923076923076923077 - 21 = 5r 51.923076923076923076923076923077 = 5r 10.384615384615384615384615384615 = r not sure if this is right

OpenStudy (anonymous):

i got the answer of t=57/13 and r=-9/13 but i have to put it into a matrices

OpenStudy (lilg132):

lol damn i had it right the first time

OpenStudy (lilg132):

let me write it out again

OpenStudy (anonymous):

ok but i dont know how to put it into a matrix

OpenStudy (lilg132):

4t - 5r = 21 3t - 7r = 18 4t = 21 + 5r divide everything by 4 t = 5.25 + 1.25r 3(5.25 + 1.25r) - 7r = 18 15.75 + 3.75r - 7r = 18 15.75 - 18 = 7r - 3.75r -2.25 = 3.25r -2.25 / 3.25 = 3.25r / 3.25 -9/13 = r 4t - 5(-9/13) = 21 4t + 45/13 = 21 4t = 21 - 45/13 4t = 228/13 4t / 4 = (228/13)/4 t = 57/13

OpenStudy (lilg132):

ok let me work out how to put into matrix

OpenStudy (anonymous):

ok i got all that i just dont no how to make a matrix out of that

OpenStudy (lilg132):

no idea with matrixes going to have to ask someone else sorry

OpenStudy (lilg132):

im out of here bye

OpenStudy (aripotta):